atlas Reference Manual

0.2.4r1

Generated by Doxygen 1.4.6

Mon Mar 20 07:08:52 2006

Contents

1 atlas Directory Hierarchy 1 1.1 atlas Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 atlas Directory Documentation 3 2.1 sources/error/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 sources/gkmod/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 sources/interface/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.4 sources/io/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.5 sources/memory/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.6 sources/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.7 sources/structure/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.8 sources/test/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.9 sources/utilities/ Directory Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Chapter 1

atlas Directory Hierarchy

1.1 atlas Directories

This directory hierarchy is sorted roughly, but not completely, alphabetically:

sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gkmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 io . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

2  atlas Directory Hierarchy

Chapter 2

atlas Directory Documentation

2.1 sources/error/ Directory Reference

Files

• file error.cpp

• file error.h

• file kl_error.cpp

• file kl_error.h

2.2 sources/gkmod/ Directory Reference

Files

• file blocks.cpp

• file blocks.h

• file blocks_fwd.h

• file bruhat.cpp

• file bruhat.h

• file bruhat_fwd.h

• file descents.h

• file descents_fwd.h

• file involutions.cpp

• file involutions.h

• file involutions_fwd.h

• file kgb.cpp

• file kgb.h

• file kgb_fwd.h

• file kl.cpp

• file kl.h

• file kl_fwd.h

• file klcomputations.cpp

• file klcomputations.h

• file klcomputations_fwd.h

• file klsupport.cpp

• file klsupport.h

• file klsupport_fwd.h

• file wgraph.cpp

• file wgraph.h

2.3 sources/interface/ Directory Reference

Files

• file commands.cpp

• file commands.h

• file commands_fwd.h

• file emptyhelp.cpp

• file emptyhelp.h

• file emptymode.cpp

• file emptymode.h

• file helpmode.cpp

• file helpmode.h

• file input.cpp

• file input.h

• file input_noreadline.c

• file input_readline.c

• file io.cpp

• file io.h

• file main.cpp

• file mainhelp.cpp

• file mainhelp.h

• file mainmode.cpp

• file mainmode.h

• file realhelp.cpp

• file realhelp.h

• file realmode.cpp

• file realmode.h

• file special.cpp

• file special.h

• file version.h

2.4 sources/io/ Directory Reference

Files

• file basic_io.cpp

• file basic_io.h

• file basic_io_def.h

• file block_io.cpp

• file block_io.h

• file cartan_io.cpp

• file cartan_io.h

• file complexredgp_io.cpp

• file complexredgp_io.h

• file complexredgp_io_fwd.h

• file interactive.cpp

• file interactive.h

• file interactive_fwd.h

• file interactive_lattice.cpp

• file interactive_lattice.h

• file interactive_lietype.cpp

• file interactive_lietype.h

• file ioutils.cpp

• file ioutils.h

• file kgb_io.cpp

• file kgb_io.h

• file kl_io.cpp

• file kl_io.h

• file layout.h

• file poset_io.cpp

• file poset_io.h

• file prettyprint.cpp

• file prettyprint.h

• file prettyprint_def.h

• file realform_io.cpp

• file realform_io.h

• file realredgp_io.cpp

• file realredgp_io.h

• file realredgp_io_fwd.h

• file realweyl.cpp

• file realweyl.h

• file realweyl_fwd.h

• file realweyl_io.cpp

• file realweyl_io.h

• file wgraph_io.cpp

• file wgraph_io.h

2.5 sources/memory/ Directory Reference

Files

• file allocator.cpp

• file allocator.h

• file allocator_def.h

• file pool.cpp

• file pool.h

• file stlset.h

• file stlvector.h

• file typenumber.cpp

• file typenumber.h

• file typestring.cpp

• file typestring.h

2.6 sources/ Directory Reference

Directories

• directory error

• directory gkmod

• directory interface

• directory io

• directory memory

• directory structure

• directory test

• directory utilities

2.7 sources/structure/ Directory Reference

Files

• file bitvector.h

• file bitvector_def.h

• file bitvector_fwd.h

• file cartan.cpp

• file cartan.h

• file cartan_fwd.h

• file cartanclass.cpp

• file cartanclass.h

• file cartanclass_fwd.h

• file complexredgp.cpp

• file complexredgp.h

• file complexredgp_fwd.h

• file dynkin.cpp

• file dynkin.h

• file gradings.cpp

• file gradings.h

• file gradings_fwd.h

• file lattice.cpp

• file lattice.h

• file lattice_def.h

• file latticetypes.cpp

• file latticetypes.h

• file latticetypes_fwd.h

• file lietype.cpp

• file lietype.h

• file lietype_fwd.h

• file prerootdata.cpp

• file prerootdata.h

• file prerootdata_fwd.h

• file realform.h

• file realredgp.cpp

• file realredgp.h

• file realredgp_fwd.h

• file rootdata.cpp

• file rootdata.h

• file rootdata_def.h

• file rootdata_fwd.h

• file subquotient.h

• file subquotient_def.h

• file subquotient_fwd.h

• file tits.cpp

• file tits.h

• file tits_fwd.h

• file topology.cpp

• file topology.h

• file tori.cpp • file tori.h

• file tori_fwd.h

• file weyl.cpp

• file weyl.h

• file weyl_fwd.h

• file weylsize.cpp

• file weylsize.h

2.8 sources/test/ Directory Reference

Files

• file kltest.cpp

• file kltest.h

• file test.cpp

• file test.h

• file testprint.cpp

• file testprint.h

• file testrun.cpp

• file testrun.h

2.9 sources/utilities/ Directory Reference

Files

• file abelian.cpp

• file abelian.h

• file abelian2_fwd.h

• file abelian_fwd.h

• file arithmetic.cpp

• file arithmetic.h

• file arithmetic_def.h

• file bitmap.cpp

• file bitmap.h

• file bitmap_def.h

• file bitmap_fwd.h

• file bits.cpp

• file bits.h

• file bitset.cpp

• file bitset.h

• file bitset_def.h

• file bitset_fwd.h

• file comparison.h

• file constants.cpp

• file constants.h

• file ctr_iterator.h

• file graph.cpp

• file graph.h

• file intutils.h

• file intutils_def.h

• file matrix.h

• file matrix_def.h

• file matrix_fwd.h

• file partition.cpp

• file partition.h

• file partition_def.h

• file polynomials.h

• file polynomials_def.h

• file polynomials_fwd.h

• file poset.cpp

• file poset.h

• file poset_fwd.h

• file set.h

• file setutils.cpp

• file setutils.h

• file setutils_def.h

• file size.cpp

• file size.h

• file size_def.h

• file smithnormal.h

• file smithnormal_def.h

• file tags.h

Index

sources/ Directory Reference, 8 sources/error/ Directory Reference, 3 sources/gkmod/ Directory Reference, 4 sources/interface/ Directory Reference, 5 sources/io/ Directory Reference, 6 sources/memory/ Directory Reference, 7 sources/structure/ Directory Reference, 9 sources/test/ Directory Reference, 11 sources/utilities/ Directory Reference, 12